OpenStudy (anonymous):

Which of the following are vertical asymptotes of the function y = 2cot(3x) + 4? Check all that apply. A. x = pi/3 B. x = +/- pi/2 C. x = 2pi D. x = 0

OpenStudy (anonymous):

cot ( pi ) is undefined. what would you plug in for x in >> cot( 3x ) to get cot (pi)?

OpenStudy (anonymous):

cot (pi) = vertical retricemtote cot (0) = vert. retricemptote.

OpenStudy (anonymous):

So is the answer X=0? Or is there another correct answer?

OpenStudy (anonymous):

cot (3x) = cot (pi) >> undefined values 3x= pi x= ? that's another solution cot (3pi/2) is also undefined 3pi/2 = 3x x=? another solution cot (-3pi/2) is undefined cot(-3pi/2) = cot(3x) -3pi/2 = 3x x=? another solution

OpenStudy (anonymous):

so there are 3 right answers to your question.

OpenStudy (anonymous):

x=pi/3 x=pi/ 2/3. That doesn't seem right. Are you implying A, C, and D?

OpenStudy (anonymous):

3pi/2 = 3x divide 3 both sides pi/2 = x -3pi/2 = 3x =pi/2 = x not C, didn't mention 2pi :)

OpenStudy (anonymous):

-pi/2 = x for the 2nd solution

OpenStudy (anonymous):

so A, D, and B, yes?

OpenStudy (anonymous):

My bad. Yeah I think that makes sense.